solve for the variable
1/3(x-2)=x/4+2

Answer by tommyt3rd(5050)

1/3(x-2)=x/4+2
x/3-2/3=x/4+2
x/3-x/4=2 2/3
(4x-3x)/12=8/3
x/12=8/3
x/4=8
x=32